OpenStudy (anonymous):

A researcher tagged 40 giant turtles in 2008. In 2009, she returned and captured 50 turtles, 16 of which were tagged. In 2010, she returned again and captured 30 turtles, 8 of which were tagged. a. Use a proportion to estimate the turtle population for 2009. b. Use a proportion to estimate the turtle population for 2010.

OpenStudy (anonymous):

@wifibasedgod

OpenStudy (anonymous):

i kinda forgot proportions (:|) (._.) (|:) (.-.)

OpenStudy (anonymous):

sorey

OpenStudy (anonymous):

help @kx2bay please

OpenStudy (anonymous):

@ganeshie8 help

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

its almost 2:00

OpenStudy (anonymous):

i remember this now

OpenStudy (anonymous):

they ask to write a fraction

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/proportions.html

OpenStudy (anonymous):

50 turtles to 30 turtles translates to 5/3 (five over three) and 16 tagged turtles to 8 tagged turtles translates to 2/1 (two over one).

OpenStudy (anonymous):

can you solve from here?

ganeshie8 (ganeshie8):

look at first reply for long answer, and second reply for short : https://in.answers.yahoo.com/question/index?qid=20130517082822AANJESo

OpenStudy (campbell_st):

this is capture -recapture and is a application of ratios let p = population estimate 2009 its modelled by Population/tagged = recatpure pop/tagged \[\frac{P}{40} = \frac{50}{16}\] just solve for P

OpenStudy (campbell_st):

for 2010 use \[\frac{P}{40} = \frac{30}{8}\] again, solve for P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is \[\frac{ P }{ 40 }\] = \[\frac{ 30 }{ 8 }\] like the formula?